OpenStudy (anonymous):

A car is traveling at 50 mi/h when the brakes are fully applied, producing a constant deceleration of 44 ft/s2. What is the distance covered before the car comes to a stop? (Give your answers correct to one decimal place.)

OpenStudy (anonymous):

help me pleasssssseeeee

OpenStudy (anonymous):

i need equation

OpenStudy (anonymous):

It doesnt give me an equation.

OpenStudy (anonymous):

this is like physics. initial velocity is 50 mi/h, acceleration is -44ft/s^2, d = ?. An equation to use would be Vf^2=Vi + 2ad. Plug in values and go for it! Make sure to get the same units! (mi/h, ft/s^2).

OpenStudy (anonymous):

thanks you've been the most help so far... but we are supposed to use the calculus way... but physics seems much more logical

OpenStudy (anonymous):

Yea, the calculus way is using derivatives I believe. Acceleration is the derivative of velocity, and velocity is the derivative of position (distance). maybe that will help... :)

OpenStudy (anonymous):

but

OpenStudy (anonymous):

using this gives me 0 = 50 * 2 * -44 * d

OpenStudy (anonymous):

ok so what do you use as your acceleration function

OpenStudy (anonymous):

We can use the equation v^2 / 2 = ad. But first we have to change mph to fps. 50*5280/3600 = 73.333 Now 73.333^2/2 = 2688.9 and dividing by 44 gives 61.1 ft for the stopping distance.

OpenStudy (anonymous):

she needs to us calculus

OpenStudy (anonymous):

exactly...

OpenStudy (anonymous):

integrals....

OpenStudy (anonymous):

well a(t)=-44

OpenStudy (anonymous):

v(t)=-44t+c 50=-44(0)+c c=50 v(t)=-44t+50 s(t)=-22t^2+50t+c

OpenStudy (anonymous):

He's on the right track

OpenStudy (anonymous):

at t=0,s(t)=0 0=-22(0)^2+50(0)+c 0=c s(t)=-22t^2+50t when the car is at a complete stop v(t)=0 0=-44t+50 -50=-44t t=25/22 s(25/22)=-22(25/22)^2+50(25/22)

OpenStudy (anonymous):

i got 28.40909091

OpenStudy (anonymous):

no thats not right.... its 61.1 but i feel like your on the right track...

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

oh i think i know what it is... the units need to be converted >_<

OpenStudy (anonymous):

does it ask for the distance in a specific? i'm guessing it wants it in meters

OpenStudy (anonymous):

the velocity needs to be switched to ft/s^2

OpenStudy (anonymous):

yep i got 61.056

OpenStudy (anonymous):

ok yes...!!!!! FANTASTIC!!!

OpenStudy (anonymous):

so using the conversion that feeman did your v(0)=73.3ft/s

OpenStudy (anonymous):

and the way you did it i've seen before and understnd.

OpenStudy (anonymous):

if you do the same step i did as above but use 73.3ft/sec instead of 50mi/hr you should get the answer

OpenStudy (anonymous):

let me know if you can't get it